OpenStudy (anonymous):

Ok, have another I can't come up with an answer that works:( 4(x+3)=3x-10. I got -22 which doesn't fit for X. Can someone help me with this one please?

OpenStudy (anonymous):

4x+12=3x-10 4x-3x+12=-10 x=-10-12 x=-22 good to me let check 4(x+3)=3x-10 4(-22+3)=3(-22)-10 4(-19)=-66-10 -76=-76 true it works im guessing you made a calculation mistake

OpenStudy (radar):

Remove parenthese 4x=12+3x-10 Subtract 3x from both sides x+12=-10 subtract 12 from both sides x=-22, why do you say it doesn't fit for x substituting -22 for x in the original equations and check it out.

OpenStudy (anonymous):

I made a mistake in checking, I must have, thank you both. I will go check again.

OpenStudy (radar):

typo alert change = sign to + in first line and change + sign to = lol.

OpenStudy (anonymous):

ok;) will do;)

OpenStudy (anonymous):

I got the 4(-22+3) part wrong, thanks you guys, I didn't see it. It's -19 and I was not doing inside ( ) first:(